OpenStudy (anonymous):

Create your own polynomial with a degree of 3 that when divided by x + 2 has a remainder of -4

myininaya (myininaya):

what do you have so far?

OpenStudy (anonymous):

Nothing, this is to help someone out and this is all she wrote me. I don't know anything about this yet.

myininaya (myininaya):

Well I will tell you some things to start you off Say we called the polynomial f(x) We are given f(-2)=-4

myininaya (myininaya):

You need at least two terms.

myininaya (myininaya):

actually you only need one term

myininaya (myininaya):

or at least one term anyways

myininaya (myininaya):

but whatever the polynomial is you must have x^3 somewhere up in there

myininaya (myininaya):

f(x)=ax^3 <-- so you can choose this

myininaya (myininaya):

f(x)=x^3+b <--this

myininaya (myininaya):

and there are a lot more varieties you can choose

myininaya (myininaya):

but you must apply the f(-2)=-4

OpenStudy (anonymous):

How does f(-2)=-4 get brought into f(x)=ax^3?

myininaya (myininaya):

f(-2)=-4 means replace x with -2 and when the x=-2 the expression =-4

myininaya (myininaya):

so you can find what that little value a is

OpenStudy (anonymous):

The expression being f(x)=ax^3?

myininaya (myininaya):

the expression ax^3 was named f(x)

OpenStudy (anonymous):

How would you find a?

myininaya (myininaya):

by using f(-2)=-4

myininaya (myininaya):

this was given in the problem

OpenStudy (anonymous):

What do I do with f(-2)=-4, then?

myininaya (myininaya):

you apply it to your function we called f

myininaya (myininaya):

remember f(x)=ax^3

myininaya (myininaya):

use f(-2)=-4 to find a

myininaya (myininaya):

f(-2)=-4 says to replace x with -2 and at x=-2 your expression=-4

OpenStudy (anonymous):

Your expression being f(x)=ax^3?

myininaya (myininaya):

pretend like we have g(x)=9x+c and I wanted to find c and I was given g(3)=4 g(3)=4 means we have g(3)=9(3)+c=4 so we have 9(3)+c=4 so 27+c=4 then c=4-27=-23 so g(x)=9x-23

myininaya (myininaya):

g(3)=4 means when x=3 the expression we called g is 4 replace x with 3 the expression at x=3, is 4 so we have 9(3)+c=4 ---

myininaya (myininaya):

you need to apply f(-2)=-4 to f(x)=ax^3 to find a

OpenStudy (anonymous):

So it'd be f(-2)=a(-2)^3=-4?

myininaya (myininaya):

yep now solve for a

OpenStudy (anonymous):

-8a=-4, divide the -8 a=0.5 ?

myininaya (myininaya):

yep you have found a function that works

myininaya (myininaya):

f(x)=.5x^3 or x^3/2 since 1/2=.5

OpenStudy (anonymous):

Yeah. How'd you get the 2nd function?

OpenStudy (anonymous):

f(x)=ax^3 <- that one

OpenStudy (anonymous):

f(x)=x^3+b <- or this one